OpenStudy (nuccioreggie):

Use absolute value to express the distance between −10 and 16 on the number line. |−10 − 16| = −6 |−10 − 16| = 6 |−10 − 16| = −26 |−10 − 16| = 26

OpenStudy (anonymous):

I-10-16I = I-26I = 26

OpenStudy (nuccioreggie):

so c

OpenStudy (nuccioreggie):

@BPDlkeme234